An exam given typically has a mean of 75 and a standard deviation of 5 points. If Tim receives a curved score of 80, about what percent of the class has a worse score than Tim?
Answer #1

\\ \mu=75, \sigma=5\\\\ Let~X~represent the score.\\\\ P(X<80)=P(\frac{X-\mu}{\sigma}<\frac{80-75}{5})\\\\ ~~~~~~~=P(Z<1)\\\\ ~~~~~~~~=0.8413~(use~z~table)\\\\ ~~~~~~~~~~=80.13\%